# Close-in compact super-Earth systems emerging from resonant chains: slow destabilization by unseen remnants of formation

### v1

@@ Acknowledgements.
First, our re-integrations varied the integrator timestep as , where ranged from and is 1/20 of the period of the innermost planet. Fig. 18 shows time to instability in these runs (solid circles). In each case, an instability of the same nature arose in the system, with the time to instability within a factor of two of the original value of 110 Myr. The absence of a trend, especially the lack of an increase in the instability timescale for small timesteps, suggests that the slow growth in libration amplitude is not caused by truncation of the N-body Hamiltonian. If this effect were purely due to roundoff error or truncation error, there would be a clear increasing or decreasing trend with the timestep, respectively.
Second, we re-integrated the outer system with the SABA(10,6,4) symplectic integrator implemented in rebound (Laskar2001; Rein2019a), which has a much higher order than WHFast. SABA indeed conserves angular momentum and energy much better than WHFast for the same timestep (thick lines in Fig. 19), at least until a close encounter that it was not designed to handle. However, the actual time of instability is nearly the same with SABA (crosses in Fig. 18) and TRACE with WHFast (circles). This suggests that that the instability is robust to choice of integrators and likely not a numerical artifact.

### v2

@@ Appendix B Numerical tests
First, our re-integrations varied the integrator timestep as , where ranged from and is 1/20 of the period of the innermost planet. Figure 18 shows time to instability in these runs (solid circles). In each case, an instability of the same nature arose in the system, with the time to instability within a factor of two of the original value of 110 Myr. Such scatter represents a fundamental limitation in instability time predictions for chaotic systems (Hussain and Tamayo 2020). The absence of a trend, especially the lack of an increase in the instability timescale for small timesteps, suggests that the slow growth in libration amplitude is not caused by truncation of the N-body Hamiltonian. If this effect were purely due to roundoff error or truncation error, there would be a clear increasing or decreasing trend with the timestep, respectively.
Second, we re-integrated the outer system with the SABA(10,6,4) symplectic integrator implemented in rebound (Laskar and Robutel 2001; Rein et al. 2019b), which has a much higher order than WHFast. SABA indeed conserves angular momentum and energy much better than WHFast for the same timestep (thick lines in Fig. 19), at least until a close encounter that it was not designed to handle. However, the actual time of instability is nearly the same with SABA (crosses in Fig. 18) and TRACE with WHFast (circles). This suggests that that the instability is robust to choice of integrators and likely not a numerical artifact.